# Disorder Suppression of Charge Density Wave in the Honeycomb Holstein Model

### v2

@@ II Model and method
In order to investigate the lattice size effect, the data are presented for $L=4,5,6$, where $L$ represents the lattice size, and for the honeycomb lattice, the number of lattice points is $N=2L^{2}$. A dimensionless electron-phonon coupling strength $\lambda_{D}=\lambda^{2}/(\omega_{0}^{2}W)$ is employed, where $W=6t$ denotes the bandwidth. To ensure that the systematic Trotter error of the system is smaller than the statistical error of Monte Carlo sampling, the discretization mesh is set as $\Delta\tau=1/20$. In the Holstein model, $\mu=-\lambda^{2}/\omega_{0}^{2}$ corresponds to the half-filling case. We set the parameter $\mu$ to maintain the charge density $\rho$ in a half-filled state $\rho=\langle\hat{n}_{\mathbf{i}}\rangle=1$. All data points are the average of 20 independent disorder configurations; the error bars correspond to the standard deviation of these 20 datasets, which comprehensively reflects both Monte Carlo sampling fluctuations and disorder configurational fluctuations[39].
@@ III.4 Electron kinetic energy, DC conductivity and spectral function
To reveal the microscopic regulatory mechanism of disorder on CDW order, we calculated the spectral function $A(\omega)$, with the results shown in Fig. 7. The system parameters were fixed as follows: lattice size $L=6$, dimensionless electron-phonon coupling $\lambda_{D}=2/3$, phonon frequency ratio $\omega_{0}/t=1$, and half-filling $\rho=1$. Fig. 7(a) corresponds to a higher temperature, where thermal fluctuations dominate the electronic behavior. The $A(\omega)$ curves exhibit broadened and dispersed features with no significant differences across different disorder strengths. This indicates that at high temperatures, the effect of thermal fluctuations far outweighs that of disorder, and the modulating effect of disorder on the spectral function is masked. Fig. 7(b) corresponds to a lower temperature, where the role of disorder becomes prominent as thermal fluctuations weaken: a strict CDW gap exists in the clean system; as $\Delta$ increases to 0.5, 1.0, and 1.5, the gap gradually becomes shallower and blurred, and the density of states at $\omega=0$ rises from strictly zero to a small finite value. This confirms that disorder disrupts long-range CDW order through Anderson localization, and even though the magnitude of the density of states remains small, it signifies the loss of the gap’s strictness. These results are consistent with the conclusions from the CDW structure factor and electron kinetic energy: at high temperatures, the effect of disorder is masked, while at low temperatures, the synergistic competition between electron-phonon coupling and Anderson localization dominates the evolution of the gap, providing microscopic support for the regulation of CDW order by disorder.

### v3

# Disorder Suppression of Charge Density Waves in the Honeycomb Holstein Model

@@ II Model and method
In order to investigate the lattice size effect, the data are presented for $L=4,5,6$, where $L$ represents the lattice size, and for the honeycomb lattice, the number of lattice points is $N=2L^{2}$. A dimensionless electron-phonon coupling strength $\lambda_{D}=\lambda^{2}/(\omega_{0}^{2}W)$ is employed, where $W=6t$ denotes the bandwidth. To ensure that the systematic Trotter error of the system is smaller than the statistical error of Monte Carlo sampling, the discretization mesh is set as $\Delta\tau=1/20$. In the Holstein model, $\mu=-\lambda^{2}/\omega_{0}^{2}$ corresponds to the half-filling case. We set the parameter $\mu$ to maintain the charge density $\rho$ in a half-filled state $\rho=\langle\hat{n}_{\mathbf{i}}\rangle=1$. We treat disorder as quenched (static). Independent static realizations of the random hopping configuration are generated. This models a frozen disorder potential and is the standard approach for studying static disorder effects, in contrast to treating disorder as a dynamical (annealed) field. For each realization, observables are computed across all temperatures using the same fixed configuration, isolating the temperature dependence. All data points are the average of 20 independent disorder configurations; the error bars correspond to the standard deviation of these 20 datasets, which comprehensively reflects both Monte Carlo sampling fluctuations and disorder configurational fluctuations[39].
@@ III.4 Electron kinetic energy, DC conductivity and spectral function
To reveal the microscopic regulatory mechanism of disorder on CDW order, we calculated the spectral function $A(\omega)$, with the results shown in Fig. 7. The system parameters were fixed as follows: lattice size $L=6$, dimensionless electron-phonon coupling $\lambda_{D}=2/3$, phonon frequency ratio $\omega_{0}/t=1$, and half-filling $\rho=1$. Fig. 7(a) corresponds to a higher temperature, where thermal fluctuations dominate the electronic behavior. The $A(\omega)$ curves exhibit broadened and dispersed features with no significant differences across different disorder strengths. This indicates that at high temperatures, the effect of thermal fluctuations far outweighs that of disorder, and the modulating effect of disorder on the spectral function is masked. Fig. 7(b) corresponds to a lower temperature, where the role of disorder becomes prominent as thermal fluctuations weaken: a strict CDW gap exists in the clean system; as $\Delta$ increases to 0.5, 1.0, and 1.5, the gap is progressively broadened and suppressed. This significant broadening of the gap edge and the suppression of the coherence peaks confirm that disorder disrupts the long-range CDW order, consistent with the mechanism of Anderson localization. These results are consistent with the conclusions from the CDW structure factor and electron kinetic energy: at high temperatures, the effect of disorder is masked, while at low temperatures, the competition between electron-phonon coupling and disorder dominates the gap’s degradation, providing microscopic support for the suppression of CDW order by disorder.